OpenStudy (anonymous):

ln 3 − 2 ln 7 + ln 14

OpenStudy (mathmate):

Hint: According to the laws of logarithms, log(ab)=log(a)+log(b) log(a/b)=log(a)-log(b).

OpenStudy (anonymous):

ln3-ln(7^2)+ln14

OpenStudy (anonymous):

subtract sign you see is same as division sign, and addition sign is same as multilication sign.

OpenStudy (anonymous):

ln (3/49) + ln 14

OpenStudy (anonymous):

so would it be ln (3/49 * 14)?

OpenStudy (noelgreco):

Yes

OpenStudy (anonymous):

thanks